OpenStudy (anonymous):

Admission to a local wave pool costs a family $22 for 4 people. what is the unit cost? choices: $ 5.50 per person $ 8.80 per person $ 3 per person $ 10 per person

OpenStudy (anonymous):

22/4

Parth (parthkohli):

= 22/4 = 5.50 dollars per person.

OpenStudy (anonymous):

i.e. 5.50

OpenStudy (anonymous):

22/4 = 5.5 ? not 5.50..

OpenStudy (anonymous):

@Tami!&_& we can put 0 as per the opt. it doesn't matter